40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Jul 22 10:13:31 2010, Alastair, Laser, GYRO, GYRO update 
    Reply  Thu Jul 22 21:27:23 2010, Zach, Laser, GYRO, GYRO update 
Message ID: 869     Entry time: Thu Jul 22 21:27:23 2010     In reply to: 866
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: GYRO update 

Not that I see how this could have changed, but did you do the standard error signal checks? If you switched a cable here or there the RF phase will have changed, and you might need to retune it and/or flip the inverter on the bluebox. I know you know this but it can't hurt to mention.

Also, when you do get the CCW locked again, before you get the other PDH box form the 40m, you might want to check that you indeed have the right beam into/out of the AOM by just looking at the CW output on the TRANS CCD. If you see a 00 somewhere around an fmod of 47.5 MHz then you're golden; if you only see higher-order modes, you probably have the once-passed first order or something, like we did one time before.

Quote:

Things generally progressed yesterday up until the evening when they started un-progressing a bit.

We got the feedback to the slow channel on and working.  It uses the DAQ and has a filter with zero at 10hz and a pole at 10mHz.  We were able to see that with the cavity locked, the piezo signal was kept around 0v for longer timescale changes.

The second beam is now going through the AOM and we have maximised the power in the first order on the way through the first time, then added an aperture to pick off only the first order before reflecting the beam back through the AOM along the same line.  At the output of this we again pickoff the first order beam and use the power meter to maximise the power by changing the position of the mirror behind the AOM.  We are getting an efficiency of approx 56% from the AOM at the moment.

We then got this beam into the cavity and got it aligned well enough to see fringes and something that looked like a 00 mode flashing.  Not well aligned, but nearly there.  At this point we connected a voltage source to the modulation input on the VCO for the AOM to scan up through the cavity in the CW direction.  We will need to collect the other PDH box from the 40m.

At this point we went back to the CCW loop to lock it - the beam had somehow got misaligned.  We spent the rest of the day trying to get the beam back into alignment with the cavity without really succeeding.  We are pretty certain that none of the cavity mirrors got moved so it should still be aligned.  We only worked on steering the beam into the cavity.  We did see 00mode resonances but it doesn't want to lock on these.  We maximised the power in transmission, and were careful to check the alignment of the beam on all the PDs but to no avail.  We will look at this again today.

 

ELOG V3.1.3-